# latexOverwriting file

### v4

@@ Appendix A Auxiliary Results
$\left\lvert\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u_{1})\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)-\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u_{2})\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)\right\rvert$
@@
$\qquad\leqslant\frac{C}{2}\left\lvert\left\lvert y-h(u_{1})\right\rvert_{\mathit{\Gamma}}^{2}-\left\lvert y-h(u_{2})\right\rvert_{\mathit{\Gamma}}^{2}\right\rvert\exp\left(-\frac{1}{3}\lvert y\rvert_{\mathit{\Gamma}}^{2}\right).$ (A.21)
Using that $a^{2}-b^{2}=(a-b)(a+b)$ and then the triangle inequality, we deduce that
$\left\lvert\left\lvert y-h(u_{1})\right\rvert_{\mathit{\Gamma}}^{2}-\left\lvert y-h(u_{2})\right\rvert_{\mathit{\Gamma}}^{2}\right\rvert$ $\leqslant\left\lvert h(u_{2})-h(u_{1})\right\rvert_{\mathit{\Gamma}}\Bigl{(}\left\lvert h(u_{1})\right\rvert_{\mathit{\Gamma}}+\left\lvert h(u_{2})\right\rvert_{\mathit{\Gamma}}+2\lvert y\rvert_{\mathit{\Gamma}}\Bigr{)}$
@@
$\leqslant\lVert\mathit{\Gamma}^{-\frac{1}{2}}\rVert\ell_{h}\lvert u_{2}-u_{1}\rvert\Bigl{(}\left\lvert h(u_{1})\right\rvert_{\mathit{\Gamma}}+\left\lvert h(u_{2})\right\rvert_{\mathit{\Gamma}}+2\lvert y\rvert_{\mathit{\Gamma}}\Bigr{)}.$ (A.22)
Here $\ell_{h}$ is the Lipschitz constant of $h$; see Assumption H. Trading a bit of the exponential decay in (A.21) in order to control the growth from the term $2\lvert y\rvert_{\mathit{\Gamma}}$ in (A.22), we obtain (A.19).
@@
In order to conclude the proof, using the definition of $\mathsf{P}$, we calculate that
@@
Appendix B Technical Results for Theorems 4 and 8
We show moment bounds in Subsection B.1, and we prove that $\mathsf{T}_{j}\mu=\mathsf{B}_{j}\mu$ for any Gaussian probability measure $\mu$ in Subsection B.2. Finally, we prove the stability results used in the proofs of Theorems 4 and 8 in Subsections B.3 and B.4, respectively.
@@ B.1 Moment Bounds
Let $\mu$ denote a probability measure on ${}^{d}$. Under Assumption H, it holds that
@@
Let $\mu$ denote a probability measure on ${}^{d}$. Under Assumption H, it holds that
@@
$\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)$ $\preccurlyeq\int_{{}^{d}}\begin{pmatrix}u\\ h(u)\end{pmatrix}\otimes\begin{pmatrix}u\\ h(u)\end{pmatrix}\,\mathsf{P}\mu(u)\,\mathrm{d}u$ (B.14)
@@
$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}$ $\geqslant(1-\varepsilon)a^{\mathsf{T}}\mathit{\Sigma}a-\left(\frac{1}{\varepsilon}-1\right)\kappa_{h}^{2}\left\lvert b\right\rvert^{2}+b^{\mathsf{T}}\mathit{\Gamma}b$ (B.24)
@@
where the supremum is over the unit sphere in ${}^{n}$, centered at the origin and in the Euclidean distance. Similarly
@@ B.3 Stability Results
$\forall j\in,\qquad d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant L_{\mathsf{T}}\,d_{g}(\pi,p),\qquad p:=\mathsf{Q}\mathsf{P}\mu.$
@@ B.4 Additional Stability Results for the Gaussian Projected Filter
Next, we show the local Lipschitz continuity of the Gaussian projection map $\mathsf{G}$. To this end, we begin by proving the following result. (The proof uses an approach similar to that employed in [16] for obtaining an upper bound on the distance between Gaussians in the usual total variation metric.)
@@
We assume from now on that $20<\lambda_{i}<$ 2 for all $i\in$. Employing the same reasoning as in [2, Lemma 3.1], we can prove
@@
$d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})\leqslant L_{\mathsf{G}}(R,n)\,d_{g}(\mu_{1},\mu_{2}),$
@@
Lemmas 21 and 22 on the Lipschitz continuity of $\mathsf{P}$ and $\mathsf{Q}$ do not depend on the specific choice of the function $g$ in Definition 1; they hold true also for the usual total variation distance. In contrast, in the proof of Lemma 26 the specific choice of $g$ is used in order to control differences between moments by means of $d_{g}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$.
@@
One may wonder whether the non-unit Lipschitz constant in Lemma 26 is inherent or just an artifact of the method of proof. A simple example shows that the Gaussian projection operator $\mathsf{G}$ is not globally Lipschitz continuous. Specifically, consider the probability distributions $\mu_{1}=\mathcal{N}(0,\varepsilon)$ with a variance $\varepsilon\leqslant 1/2$ and
$\mu_{2}=\varepsilon\delta_{-1}+(1-2\varepsilon)\mu_{1}+\varepsilon\delta_{1}.$
@@
$\frac{d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})}{d_{g}(\mu_{2},\mu_{2})}\xrightarrow[\varepsilon\to 0]{}\infty.$
@@
$\forall j\in,\qquad d_{g}\bigl{(}\mathsf{B}_{j}\pi,\mathsf{B}_{j}p\bigr{)}\leqslant L_{\mathsf{B}}(R,\kappa_{y})\,d_{g}\bigl{(}\pi,p\bigr{)}.$ (B.48)
@@
Appendix C Technical Results for Corollaries 6 and 9
@@
[…]
@@
$\Delta_{n}(u,v):=\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)-\exp\left(-\frac{1}{2}\left\lvert u-\Psi_{n}(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)$
Multiplying and dividing by the density of $\mathcal{N}(0,2\mathit{\Sigma})$ in the outermost integral, we obtain that
$d_{g}(\mathsf{P}\mu,\mathsf{P}_{n}\mu)$ $\leqslant\frac{1}{\sqrt{(2\pi)^{d}\det\mathit{\Sigma}}}\int_{{}^{d}}\int_{{}^{d}}\exp\left(\frac{1}{2}\left\lvert u\right\rvert_{2\mathit{\Sigma}}^{2}\right)g(u)\left|\Delta_{n}(u,v)\right|\,\mu(\mathrm{d}v)\,\exp\left(-\frac{1}{2}\left\lvert u\right\rvert_{2\mathit{\Sigma}}^{2}\right)\,\mathrm{d}u$
$\leqslant\sqrt{2^{d}}\sup_{u,v\in^{d}}\left\{\exp\left(\frac{1}{2}\left\lvert u\right\rvert_{2\mathit{\Sigma}}^{2}\right)g(u)\bigl{\lvert}\Delta_{n}(u,v)\bigr{\rvert}\right\}.$ (C.2)
In order to prove the first limit in (C.1), it remains to show that the supremum in (C.2) tends to 0 in the limit as $n\to\infty$. To this end, suppose for contradiction that there was a sequence $(u_{n},v_{n})\in^{d}\times^{d}$ and $\alpha>0$ such that
$\forall n\in\mathbf{N},\qquad\exp\left(\frac{1}{2}\left\lvert u_{n}\right\rvert_{2\mathit{\Sigma}}^{2}\right)g(u_{n})\bigl{\lvert}\Delta_{n}(u_{n},v_{n})\bigr{\rvert}\geqslant\alpha.$ (C.3)
By the assumption that $\Psi_{n}\to\Psi$ in $L^{\infty}(^{d})$, it holds that $\Delta_{n}(u_{n},v_{n})\to 0$, and so it must necessarily hold that $\lvert u_{n}\rvert\to\infty$ for (C.3) to hold. But this leads to a contradiction because, by Young’s inequality, it holds for all $\delta>0$ that
$\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\geqslant\frac{1}{1+\delta}\left\lvert u\right\rvert_{\mathit{\Sigma}}^{2}-\frac{1}{\delta}\left\lvert\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}.$
Therefore, given that $\Psi_{n}\to\Psi$ in $L^{\infty}(^{d})$, the term $\Delta_{n}(u_{n},v_{n})$ in (C.3) decreases to 0 in the limit $\lvert u_{n}\rvert\to\infty$ faster than the other terms grow.
@@
$\widetilde{\Delta}_{n}(u,y):=\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)-\exp\left(-\frac{1}{2}\bigl{\lvert}y-h_{n}(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)$
The innermost integral in (C.4), which is independent of $\mu$, converges to 0 uniformly in $u\in^{d}$, and so the conclusion follows. ∎
@@
Let $\mathsf{P}_{n}$ and $\mathsf{Q}_{n}$ denote the maps (1.5) and (1.7) (associated with the functions $\Psi_{n}$ and $h_{n}$) and assume that they too satisfy Assumption H. Denote by $(\mu^{n}_{j})_{j\in}$ the true filtering distribution associated with functions $\Psi_{n},h_{n}$. Then, with the same notation as in Lemma 30, it holds that
$\lim_{n\to\infty}\max_{j\in}d_{g}(\mathsf{G}\mathsf{Q}_{n}\mathsf{P}_{n}\mu^{n}_{j},\mathsf{Q}_{n}\mathsf{P}_{n}\mu^{n}_{j})=0.$

### v6

@@ Appendix A Auxiliary Results
$\quad+C\lvert u_{1}-u_{2}\rvert\int_{{}^{d}}\exp\left(-\frac{1}{3}\left\lvert u_{2}-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)\,\mu(\mathrm{d}v).$
By Young’s inequality, it holds for all $\delta>0$ that
@@
$\forall(a,b)\in^{d}\times^{d},\qquad\left\lvert a-b\right\rvert_{\mathit{\Sigma}}^{2}\geqslant\frac{1}{1+\delta}\left\lvert a\right\rvert_{\mathit{\Sigma}}^{2}-\frac{1}{\delta}\left\lvert b\right\rvert_{\mathit{\Sigma}}^{2}.$ (A.14)
Using this inequality with $\delta=\frac{1}{3}$ together with the assumption that $\Psi$ is bounded, we deduce that
@@
$\lvert\mathsf{P}\mu(u_{1})-\mathsf{P}\mu(u_{2})\rvert$ $\leqslant C\lvert u_{1}-u_{2}\rvert\int_{{}^{d}}\exp\left(-\frac{1}{4}\left\lvert u_{1}\right\rvert_{\mathit{\Sigma}}^{2}\right)\,\mu(\mathrm{d}v)+C\lvert u_{1}-u_{2}\rvert\int_{{}^{d}}\exp\left(-\frac{1}{4}\left\lvert u_{2}\right\rvert_{\mathit{\Sigma}}^{2}\right)\,\mu(\mathrm{d}v)$
$\leqslant 2C\lvert u_{1}-u_{2}\rvert\exp\left(-\frac{1}{4}\min\Bigl{\{}\lvert u_{1}\rvert_{\mathit{\Sigma}}^{2},\lvert u_{2}\rvert_{\mathit{\Sigma}}^{2}\Bigr{\}}\right).$ (A.15)
Next, letting $h_{s}=(1-s)h(u_{1})+sh(u_{2})$ and using a reasoning similar to that in the proof of Lemma 13, we obtain the following inequalities, which hold for all $(u_{1},u_{2},y)\in^{d}\times^{d}\times^{K}$:
$\Bigl{\lvert}\mathcal{N}\bigl{(}h(u_{1}),\mathit{\Gamma}\bigr{)}(y)-\mathcal{N}\bigl{(}h(u_{2}),\mathit{\Gamma}\bigr{)}(y)\Bigr{\rvert}$ $\leqslant\lvert h(u_{2})-h(u_{1})\rvert_{\mathit{\Gamma}}\int_{0}^{1}\lvert h_{s}-y\rvert_{\mathit{\Gamma}}\exp\left(-\frac{1}{2}\left\lvert h_{s}-y\right\rvert_{\mathit{\Gamma}}^{2}\right)\,\mathrm{d}s$
$\leqslant C\lvert u_{2}-u_{1}\rvert\exp\left(-\frac{1}{4}\lvert y\rvert_{\mathit{\Gamma}}^{2}\right),$ (A.16)
@@
where we used the Lipschitz continuity of $h$, together with (A.14) and the boundedness of $h$, in the last inequality. In order to conclude the proof, using the definition of $\mathsf{P}$, we calculate that
@@
Appendix B Technical Results for Theorems 4 and 9
We show moment bounds in Subsection B.1, and we prove that $\mathsf{T}_{j}\mu=\mathsf{B}_{j}\mu$ for any Gaussian probability measure $\mu$ in Subsection B.2. Finally, we prove the stability results used in the proofs of Theorems 4 and 9 in Subsections B.3 and B.4, respectively.
@@ B.1 Moment Bounds
Let $\mu$ denote a probability measure on d. Under Assumption H, it holds that
@@
Let $\mu$ denote a probability measure on d. Under Assumption H, it holds that
@@
$\mathcal{C}(\phi_{\sharp}\mathsf{P}\mu)$ $\preccurlyeq\int_{{}^{d}}\begin{pmatrix}u\\ h(u)\end{pmatrix}\otimes\begin{pmatrix}u\\ h(u)\end{pmatrix}\,\mathsf{P}\mu(u)\,\mathrm{d}u$
@@
$\begin{pmatrix}a\\ b\end{pmatrix}^{\mathsf{T}}\mathcal{C}(\mathsf{Q}\mathsf{P}\mu)\begin{pmatrix}a\\ b\end{pmatrix}$ $\geqslant(1-\varepsilon)a^{\mathsf{T}}\mathit{\Sigma}a-\left(\frac{1}{\varepsilon}-1\right)\kappa_{h}^{2}\left\lvert b\right\rvert^{2}+b^{\mathsf{T}}\mathit{\Gamma}b$
@@
where the supremum is over the unit sphere in n, centered at the origin and in the Euclidean distance. Similarly
@@ B.3 Stability Results
$\forall j\in\llbracket 1,J\rrbracket,\qquad d_{g}(\mathsf{T}_{j}\pi,\mathsf{T}_{j}p)\leqslant L_{\mathsf{T}}\,d_{g}(\pi,p),\qquad p:=\mathsf{Q}\mathsf{P}\mu.$
@@ B.4 Additional Stability Results for the Gaussian Projected Filter
Next, we show the local Lipschitz continuity of the Gaussian projection map $\mathsf{G}$. To this end, we begin by proving the following result. (The proof uses an approach similar to that employed in [2018arXiv181008693D] for obtaining an upper bound on the distance between Gaussians in the usual total variation metric.)
@@
We assume from now on that $20<\lambda_{i}<$ 2 for all $i\in\llbracket 1,n\rrbracket$. Employing the same reasoning as in [MR2730330, Lemma 3.1], we can prove
@@
$d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})\leqslant L_{\mathsf{G}}(R,n)\,d_{g}(\mu_{1},\mu_{2}),$ (B.31)
@@
Lemmas 23 and 24 on the Lipschitz continuity of $\mathsf{P}$ and $\mathsf{Q}$ do not depend on the specific choice of the function $g$ in Definition 1; they hold true also for the usual total variation distance. In contrast, in the proof of Lemma 28 the specific choice of $g$ is used in order to control differences between moments by means of $d_{g}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet},\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})$.
@@
A simple example shows that the constant $L_{\mathsf{G}}(R,n)$ in (B.31) is divergent in the limit $R\to\infty$, indicating that the Gaussian projection operator $\mathsf{G}$ is not globally Lipschitz continuous. Specifically, take $\varepsilon\leqslant\frac{1}{2}$ and consider the probability distributions
$\mu_{1}=\mathcal{N}(0,\varepsilon),\qquad\mu_{2}=\varepsilon\delta_{-1}+(1-2\varepsilon)\mu_{1}+\varepsilon\delta_{1}.$
@@
$\frac{d_{g}(\mathsf{G}\mu_{1},\mathsf{G}\mu_{2})}{d_{g}(\mu_{1},\mu_{2})}\xrightarrow[\varepsilon\to 0]{}\infty.$
@@
$\forall j\in\llbracket 0,J-1\rrbracket,\qquad d_{g}\bigl{(}\mathsf{B}_{j}\pi,\mathsf{B}_{j}p\bigr{)}\leqslant L_{\mathsf{B}}(R,\kappa_{y})\,d_{g}\bigl{(}\pi,p\bigr{)}.$ (B.33)
@@
Appendix C Technical Results for Corollaries 7 and 10
@@
$\sup_{\mu\in\mathcal{P}(^{d})}d_{g}(\mathsf{P}\mu,\mathsf{P}_{n}\mu)\xrightarrow[n\to\infty]{}0\qquad\text{and}\qquad\sup_{\mu\in\mathcal{P}_{R}(^{d})}d_{g}(\mathsf{Q}\mu,\mathsf{Q}_{n}\mu)\xrightarrow[n\to\infty]{}0,$ (C.1)
@@
$\Delta_{n}(u,v):=\exp\left(-\frac{1}{2}\left\lvert u-\Psi(v)\right\rvert_{\mathit{\Sigma}}^{2}\right)-\exp\left(-\frac{1}{2}\left\lvert u-\Psi_{n}(v)\right\rvert_{\mathit{\Sigma}}^{2}\right).$
Letting $\psi_{s}(v)=(1-s)\Psi(v)+s\Psi_{n}(v)$ and using the same reasoning as in Lemma 13, we obtain that
$\bigl{\lvert}\Delta_{n}(u,v)\bigr{\rvert}$ $\leqslant\bigl{\lvert}\Psi(v)-\Psi_{n}(v)\bigr{\rvert}_{\mathit{\Sigma}}\int_{0}^{1}\bigl{\lvert}u-\psi_{s}\bigr{\rvert}_{\mathit{\Sigma}}\,\exp\left(-\frac{1}{2}\bigl{\lvert}u-\psi_{s}\bigr{\rvert}_{\mathit{\Sigma}}^{2}\right)\,\mathrm{d}s$
$\leqslant\bigl{\lvert}\Psi(v)-\Psi_{n}(v)\bigr{\rvert}_{\mathit{\Sigma}}\int_{0}^{1}\Bigl{(}\left\lvert u\right\rvert_{\mathit{\Sigma}}+\left\lvert\psi_{s}\right\rvert_{\mathit{\Sigma}}\Bigr{)}\,\exp\left(-\frac{1}{3}\bigl{\lvert}u\bigr{\rvert}_{\mathit{\Sigma}}^{2}+\bigl{\lvert}\psi_{s}\bigr{\rvert}_{\mathit{\Sigma}}^{2}\right)\,\mathrm{d}s$ (C.2a)
$\leqslant C\bigl{\lvert}\Psi(v)-\Psi_{n}(v)\bigr{\rvert}_{\mathit{\Sigma}}\exp\left(-\frac{1}{4}\bigl{\lvert}u\bigr{\rvert}_{\mathit{\Sigma}}^{2}\right).$ (C.2b)
In (C.2a), we used that by Young’s inequality, it holds for all $\delta>0$ that
$\left\lvert u-\psi_{s}\right\rvert_{\mathit{\Sigma}}^{2}\geqslant\frac{1}{1+\delta}\left\lvert u\right\rvert_{\mathit{\Sigma}}^{2}-\frac{1}{\delta}\left\lvert\psi_{s}\right\rvert_{\mathit{\Sigma}}^{2}.$
Then in (C.2b), we used the bound $\lVert\psi_{s}\rVert_{L^{\infty}}\leqslant\kappa_{\Psi}$, which holds for all $s\in[0,1]$ by Assumption (H2). The first limit in (C.1) then follows immediately.
@@
$\widetilde{\Delta}_{n}(u,y):=\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right)-\exp\left(-\frac{1}{2}\bigl{\lvert}y-h_{n}(u)\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right).$
Using the same strategy as above, we obtain that
$\forall(u,y)\in^{d}\times^{K},\qquad\bigl{\lvert}\widetilde{\Delta}_{n}(u,y)\bigr{\rvert}\leqslant C\bigl{\lvert}h(u)-h_{n}(u)\bigr{\rvert}_{\mathit{\Gamma}}\exp\left(-\frac{1}{4}\bigl{\lvert}y\bigr{\rvert}_{\mathit{\Gamma}}^{2}\right).$
Substituting in (C.3) gives the second limit in (C.1). ∎
@@
Let $\mathsf{P}_{n}$ and $\mathsf{Q}_{n}$ denote the maps (1.5) and (1.7) (associated with the functions $\Psi_{n}$ and $h_{n}$) and assume that they too satisfy Assumption H. Denote by $(\mu^{n}_{j})_{j\in\llbracket 1,J\rrbracket}$ the true filtering distribution associated with functions $\Psi_{n},h_{n}$. Then, with the same notation as in Lemma 32, it holds that
$\lim_{n\to\infty}\max_{j\in\llbracket 0,J-1\rrbracket}d_{g}(\mathsf{G}\mathsf{Q}_{n}\mathsf{P}_{n}\mu^{n}_{j},\mathsf{Q}_{n}\mathsf{P}_{n}\mu^{n}_{j})=0.$